# Two-Stage Stance Labeling: User-Hashtag Heuristics with Graph Neural Networks thanks: Army Research Office

### v1

@@ IV Results
We conduct two sets of evaluation for method comparison: 1) cross validation over the soft-labeled users from the hashtag-labeling heuristics; and 2) human evaluation on a set of manually annotated users. With the set of manually annotated users, we also classify each user’s stance using zero-shot classification with GPT4 to evaluate the performance of state-of-the-art large language models (LLMs) as a potential alternative to manual annotation. We report the macro averaged precision, recall, and F1 scores and we use the F1 score as the primary metric due to data imbalance. We then select the best model to classify all users in the gun control and climate change datasets to obtain their stance labels. These labels are then used to conduct a case study and comparison between polarization on these topics on Twitter.
IV-A User-Hashtag Labeling
While the user-hashtag bipartite label propagation method can quickly label a number of users with high precision. It suffers from the limitation that a large proportion of users do not use hashtags, and while it is well suited for determining hashtags that are highly associated with one particular stance group as compared with the other, hashtags that are used across both stances albeit in different contexts will remain unlabeled. Table IV As such, the method has low recall across the entire user set. Despite these limitations, the method can facilitate semi-supervised training of our socially-infused text mining labeling method that leverages the textual content of user tweets and social interactions between users.
TABLE IV: Top hashtags associated with each stance group.
Climate Change
Believe Disbelieve
climatechange globalwarming
cop26 ecocredit
climate co2
actonclimate covid
climatecrisis un
climateaction russia
ipcc maga
envrionment podestaemails8
climateemergency davos2022
sustainability ukraine
Gun Control
Pro Anti
guncontrol 2a
guncontrolnow secondamendment
gunviolence guns
nra 2ndamendment
endgunviolence breaking
uvalde scotus
gunreform shallnotbeinfringed
gunreformnow selfdefense
texas usa
gop jeroneforcongress
To determine the benefits of the reciprocal hashtag labeling heuristic, we compare the performance of the GNN stance labeling models trained on the initial set of users labeled by the seed hashtags against the performance of the GNN models for predicting user stance when trained by the set of users labeled after allowing the label propagation heuristic to converge. The initial set of labeled users consists of approximately 100,000 users for the gun control dataset and 200,000 users for the climate change dataset. The set of users labeled using reciprocal label propagation consists of approximately 400,000 users for the gun control dataset and one million users for the climate change dataset. In all, the labeling heuristic resulted in a four to five fold increase in the size of the training set of labeled users available for training the GNN stance labeling models.
For the gun control dataset, training GNN models with the expanded user set from the stance labeling heuristic results in a five to ten percent increase in precision, recall, and F1 score for all three of the graph neural network architectures. By contrast, for the climate change dataset, classification performance when training with the heuristic-labeled user set declined for all GNN architectures.
IV-B Manual Labeling
We evaluate the performance of our two-stage stance labeling procedure and the benefit of including social network connectivity for stance labeling by evaluating model predictive performance against a small set of gold standard users annotated manually by domain experts. Compared with the cross-validation performance of models evaluated on the set of heuristically, performance
IV-C GPT4 Labeling
We further examine the benefit of our two-stage stance labeling by performing zero-shot classification of user stances using GPT4 for the same sample of users annotated manually. Large language models like GPT4 have shown great promise in a broad variety of natural language tasks, and we evaluate whether GPT4 can match or exceed the performance of GNN models using just textual information. We hypothesize that by incorporating the network structural information through GNNs allows models to better encode not just a user’s utterances about a topic but also there position in the network and with whom they interact; thus allowing such models to better classify user stances on particular topics.
Contrary to the results on the cross-validation, Table LABEL:tab:gpt4-results shows that GNN models outperform tesx-based models for the Climate Change dataset, while stance prediction of users in the Gun Control dataset with GNN models performed worse than text-based models.
Enriching the user representations from text-based models with social network information improves the stance labeling performance but is not able to match the classification performance of user tweets using GPT-4
Notably, for the gun control test users, twelve users were labeled with an equivalent number of tweets from both stances by GPT-4
@@ IV-C GPT4 Labeling
BERT-base-uncased 0. 0. 0. 70.35 73.02 72.52
BERT-large-uncased 0. 0. 0. 74.43 75.96 75.24
RoBERTa-base 0. 0. 0. 70.28 70.62 71.85
RoBERTa-large 0. 0. 0. 71.86 71.84 71.63
BERTweet 0 0 0 77.99 80.40 77.52
GPT-4 - - - 90.02 88.45 89.14
GraphSAGE 86.83 83.30 80.36 84.64 82.93 81.76
GAT 84.96 87.52 84.91 88.00 86.95 86.18
GIN 89.64 82.49 80.78 81.87 81.52 80.48
TABLE V: Results for stance labeling classification on the Gun Control dataset ($N=350$). The best F1 (macro) scores for each model type are shown in bold and the best overall scores are underlined.
@@
BERT-base-uncased 0. 0. 0. 63.32 66.41 63.75
BERT-large-uncased 0. 0. 0. 67.89 71.51 69.01
RoBERTa-base 0. 0. 0. 64.44 62.28 63.93
RoBERTa-large 0. 0. 0. 67.64 68.43 67.67
BERTweet 0 0 0 60.78 52.09 43.96
GPT-4 - - - 85.73 86.25 86.51
GraphSAGE 76.83 75.30 72.36 79.37 79.05 76.88
GAT 74.96 67.52 72.91 76.71 79.83 73.50
GIN 69.64 62.49 65.78 73.94 74.42 67.57
TABLE VI: Results for stance labeling classification on the Climate Change dataset ($N=250$). The best F1 (macro) scores for each model type are shown in bold and the best overall scores are underlined.
IV-D Discussion
@@
When researching the area of online polarization, there are serious divides in how the questions are structured and the areas of focus. This divide is also reflected in how different researchers integrate the literature and findings into future work. Oftentimes, this split can be seen in how social science researchers integrate computer science research into their studies and vice versa. Much of this is focused on the idea that both research groups are missing a key methodological piece the other offers. When researching online polarization, this methodological gap can inhibit the understanding of the black box of online polarization and radicalization. Big data research can provide scalability not feasible by traditional qualitative social science research and social science expertise provides the understanding of the subtleties of these interactions on human behavior. As has been seen in the aftermath of events like radicalized mass shooters, the online behavior of these individuals is not well understood and the need to understand how and why these processes take place online needs to be examined. The integration of subject matter expertise and AI algorithmic work is necessary for the large-scale application of big data methodology into human-level polarization research and policy work. These two methodologies, the qualitative expertise provided by subject matter experts, and the quantitative capabilities of computational researchers need to be combined so that the benefits of both are fully exploited. So the question becomes, “how do we balance the benefits of both methodologies to build the accuracy of our models?” This requires a balance of the nuance provided by subject matter experts with the speed and volume provided by computation experts. The ability to combine these two approaches can greatly benefit the understanding of online polarization.
The coding of blatant or straight-forward tweets to label the stance or polarization of the sentiment should be easily handled by AI algorithms. The difficulty comes in where the tweets are subtle, nuanced, or coded. It is in this gray space that qualitative analysis provides the necessary understanding to the complexities of human experiences, behaviors, and social phenomena. To accurately code this subset of tweets or users, the underlying meanings, and contexts have to be understood since these factors can impact how the tweets are meant to be interpreted. Quantitative research offers methodological tools like narrative analysis to analyze how culturally and historically contingent the terms, beliefs and issues narrators address in their writings are (Rosenwald, and Ochberg, 1992). Using tweets is different from traditional narrative analysis since they lack the more linear coherence and structure of “master” narratives and are instead characterized by brevity, terseness, and dialogism (Boje, 2001). Considering the utilization of Twitter [X] for social movements, political campaigning, and radicalization, these small narratives created from tweets can provide insights into the range of ongoing arguments and contentions that create collective identity online (Coupland and Brown, 2004). The benefit of this approach is that it allows for the analysis of culturally, historically, locationally nuanced and complex language which is exceptionally valuable when studying areas that are traditionally hidden or stealth online, such as alternative religions/cults, criminality, black markets, or non-traditional lifestyles/ communities (i.e. furries, incels, etc.). The wealth of this data adds critical information to understanding how people discuss their thoughts and opinions about polarizing topics like gun rights or climate change.
@@
One area of difficulty in the process of automating AI algorithms to detect stances on topics like gun control or climate change, is that the aim is to dichotomize beliefs that exist on a spectrum. The extreme ends of these beliefs are straight-forward to categorize (i.e. individuals who do not believe climate change is real or want to ban all guns) the majority of people exist in the in-between space. So the goal then becomes to qualitatively code the tipping point as people’s stance moves between pro-and anti-gun and believe or disbelieve climate change. It is in the middle space, where coding becomes the most difficult and language nuance becomes critical to coding schemes. An example of how automating stance analysis can mislabel users because of difficulties in discerning written cues like tone, connotation, coded words, or pictures. Some of these coding errors are due to a purposeful misdirection of a tweet that is clarified by the picture. For example, user 2945287090 said “I’m pro choice…are you? https://t.co/WtYW6zohSB” but the picture had them wearing a shirt that said pro-choice but had an array of guns to “choose” from. Since the picture cannot be assessed by the algorithm, these misdirections would be potentially miscoded. A more complex example of this is a tweet by user 18643437 “A man who live streamed himself on Facebook driving around Memphis shooting strangers was caught on camera smiling after being arrested by police. https://t.co/nSYQMPric3.” The picture is a Black man in the back of a police car smiling. The coded language around this is focused on the race of the shooter as a dog-whistle and the underlying understanding by pro-gun advocates that the race of the shooter indicates that this is not what is meant by “law-abiding gun owner.” It shifts the focus away from the issues around gun violence as an overarching problem and onto scapegoating Black males as criminals who enjoy violence. There are also users who frame their arguments in neutral language or the language of the opposite side. For example, user 66811907 states “Gun control advocates do not have a monopoly on outrage or sorrow. They do not have a monopoly on ideas for addressing gun violence. They do not have a monopoly wanting to fix the problem.” The larger context of this user’s tweets balance between being pro- Second Amendment and being critical to the responses to mass shooting events, such as Uvalde. As a whole, this user falls more strongly on the side of being anti- gun control but certainly has criticism of open market gun purchasing. The neutral framing can be difficult to parse since the full narrative of the users’ tweets need to be taken into account. These different examples highlight some of the language and contextual cues and codes that need to be contextualized.
@@
The coding errors are not only on the side of AI, however. One of the key limitations to only using human-based qualitative coding is that this type of deep understanding of stances is highly time consuming and delayed. Especially in online space, the posting moves far faster than researchers can code by hand. For mis-coded users, there are hundreds to thousands of tweets for one user. This issue means that oftentimes qualitative researchers either focus on a subset of tweets or a subset of users to be able to have insights about the online discussions of topics like guns or climate change. When examining a subset of tweets to establish stance, there will be errors. Since the coder is only looking at a section of the tweets, often randomly chosen, may not be an accurate representation of the user’s online stance. The balance between efficiency and accuracy means that there will be mis-coded information that could be “fixed” by looking at all the available data. The more narrow focus of qualitative researchers can limit the generalizability of these findings. Which for topics like polarization and radicalization, the limited generalizability then in turn limits the impact on policy and practice. The ability to integrate big-data/ AI methodology can be a force multiplier for qualitative researchers as a mixed methods approach that can improve the generalizability of these findings.
@@
It should also be noted, as was mentioned earlier, that these stances exist on a spectrum and therefore there will be individuals who could have opposite coded stances by even two human qualitative coders. Since the researcher is making a qualitative decision of when a person’s stance moves between two ends, the middle ground can often be coded either way. For example, is a person who owns guns for hunting but pushes for very strong hand gun control measures, pro gun or anti gun? Does wanting strong background checks or other restrictions inherently make someone pro/anti gun? Or does someone who has a strong stance against their personal gun ownership but also states they do not think background checks are necessary or valuable (i.e. guns aren’t for me but I don’t care about others), pro or anti gun control? Even in the context of a large body of tweets, it may not be fully clear how this stance should be coded and is not necessarily an error on either side but rather a decision reflecting this balance. An example is user 66811907 who has a large body of tweets that different coders could code differently “This your friendly reminder that New York already bans ”assault weapons,” that the shooter procured his NY-compliant ”non-assault weapon” legally, and that the cops who stopped him showed up with their ”assault weapon” version of the same gun to combat a civilian threat. https://t.co/KLmqwQM73f” and “Rifles of any kind account for fewer annual homicides than hands and feet. Mass shooters don’t need pistol grips, etc., to cause horrific carnage, but on the flip side those features can be quite useful for lawful owners using them in self-defense. That’s why cops carry them, too.” This example is just one that highlights that neither AI nor human coding is without issues, error, or disagreement. This is seen in the reporting of inter-rater reliability in qualitative coding and could be a useful measure to add to big data analysis.
The effective coding of large data sets with an aim toward integrating these two methodologies should be the goal of research aimed at informing the larger research field and policy. Without the purposeful integration of these two methodologies, much of this research, and its impact, will continue to be limited. Both guns and climate change have a range of historical contexts, political connotations and expectations, intersections with race and gender, and sometimes conspiracy theories that can play a role in how people talk about these ideas online and inform their stances. It is in this space that the subject matter expertise can provide that contextual information that can inform the work of AI algorithms. This concept has been discussed through the idea of grounding AI with use-case specific information that is outside of the training data. The social science insights provide the ground truth necessary to properly ground AI focused on complex human stances to ensure that the AI model is acknowledging the real-world context, to minimizing errors and maximizing the relevance of the generated outputs.
IV-E Discussion
When researching the area of online polarization, there are serious divides in how the questions are structured and the areas of focus. This divide is also reflected in how different researchers integrate the literature and findings into future work. Oftentimes, this split can be seen in how social science researchers integrate computer science research into their studies and vice versa. Much of this is focused on the idea that both research groups are missing a key methodological piece the other offers. When researching online polarization, this methodological gap can inhibit the understanding of the black box of online polarization and radicalization. Big data research can provide scalability not feasible by traditional qualitative social science research and social science expertise provides the understanding of the subtleties of these interactions on human behavior. As has been seen in the aftermath of events like radicalized mass shooters, the online behavior of these individuals is not well understood and the need to understand how and why these processes take place online needs to be examined. The integration of subject matter expertise and AI algorithmic work is necessary for the large-scale application of big data methodology into human-level polarization research and policy work. These two methodologies, the qualitative expertise provided by subject matter experts, and the quantitative capabilities of computational researchers need to be combined so that the benefits of both are fully exploited. So the question becomes, “how do we balance the benefits of both methodologies to build the accuracy of our models?” This requires a balance of the nuance provided by subject matter experts with the speed and volume provided by computation experts. The ability to combine these two approaches can greatly benefit the understanding of online polarization.
The coding of blatant or straight-forward tweets to label the stance or polarization of the sentiment should be easily handled by AI algorithms. The difficulty comes in where the tweets are subtle, nuanced, or coded. It is in this gray space that qualitative analysis provides the necessary understanding to the complexities of human experiences, behaviors, and social phenomena. To accurately code this subset of tweets or users, the underlying meanings, and contexts have to be understood since these factors can impact how the tweets are meant to be interpreted. Quantitative research offers methodological tools like narrative analysis to analyze how culturally and historically contingent the terms, beliefs and issues narrators address in their writings are (Rosenwald, and Ochberg, 1992). Using tweets is different from traditional narrative analysis since they lack the more linear coherence and structure of “master” narratives and are instead characterized by brevity, terseness, and dialogism (Boje, 2001). Considering the utilization of Twitter [X] for social movements, political campaigning, and radicalization, these small narratives created from tweets can provide insights into the range of ongoing arguments and contentions that create collective identity online (Coupland and Brown, 2004). The benefit of this approach is that it allows for the analysis of culturally, historically, locationally nuanced and complex language which is exceptionally valuable when studying areas that are traditionally hidden or stealth online, such as alternative religions/cults, criminality, black markets, or non-traditional lifestyles/ communities (i.e. furries, incels, etc.). The wealth of this data adds critical information to understanding how people discuss their thoughts and opinions about polarizing topics like gun rights or climate change.
One area of difficulty in the process of automating AI algorithms to detect stances on topics like gun control or climate change, is that the aim is to dichotomize beliefs that exist on a spectrum. The extreme ends of these beliefs are straight-forward to categorize (i.e. individuals who do not believe climate change is real or want to ban all guns) the majority of people exist in the in-between space. So the goal then becomes to qualitatively code the tipping point as people’s stance moves between pro-and anti-gun and believe or disbelieve climate change. It is in the middle space, where coding becomes the most difficult and language nuance becomes critical to coding schemes. An example of how automating stance analysis can mislabel users because of difficulties in discerning written cues like tone, connotation, coded words, or pictures. Some of these coding errors are due to a purposeful misdirection of a tweet that is clarified by the picture. For example, user 2945287090 said “I’m pro choice…are you? https://t.co/WtYW6zohSB” but the picture had them wearing a shirt that said pro-choice but had an array of guns to “choose” from. Since the picture cannot be assessed by the algorithm, these misdirections would be potentially miscoded. A more complex example of this is a tweet by user 18643437 “A man who live streamed himself on Facebook driving around Memphis shooting strangers was caught on camera smiling after being arrested by police. https://t.co/nSYQMPric3.” The picture is a Black man in the back of a police car smiling. The coded language around this is focused on the race of the shooter as a dog-whistle and the underlying understanding by pro-gun advocates that the race of the shooter indicates that this is not what is meant by “law-abiding gun owner.” It shifts the focus away from the issues around gun violence as an overarching problem and onto scapegoating Black males as criminals who enjoy violence. There are also users who frame their arguments in neutral language or the language of the opposite side. For example, user 66811907 states “Gun control advocates do not have a monopoly on outrage or sorrow. They do not have a monopoly on ideas for addressing gun violence. They do not have a monopoly wanting to fix the problem.” The larger context of this user’s tweets balance between being pro- Second Amendment and being critical to the responses to mass shooting events, such as Uvalde. As a whole, this user falls more strongly on the side of being anti- gun control but certainly has criticism of open market gun purchasing. The neutral framing can be difficult to parse since the full narrative of the users’ tweets need to be taken into account. These different examples highlight some of the language and contextual cues and codes that need to be contextualized.
The coding errors are not only on the side of AI, however. One of the key limitations to only using human-based qualitative coding is that this type of deep understanding of stances is highly time consuming and delayed. Especially in online space, the posting moves far faster than researchers can code by hand. For mis-coded users, there are hundreds to thousands of tweets for one user. This issue means that oftentimes qualitative researchers either focus on a subset of tweets or a subset of users to be able to have insights about the online discussions of topics like guns or climate change. When examining a subset of tweets to establish stance, there will be errors. Since the coder is only looking at a section of the tweets, often randomly chosen, may not be an accurate representation of the user’s online stance. The balance between efficiency and accuracy means that there will be mis-coded information that could be “fixed” by looking at all the available data. The more narrow focus of qualitative researchers can limit the generalizability of these findings. Which for topics like polarization and radicalization, the limited generalizability then in turn limits the impact on policy and practice. The ability to integrate big-data/ AI methodology can be a force multiplier for qualitative researchers as a mixed methods approach that can improve the generalizability of these findings.
It should also be noted, as was mentioned earlier, that these stances exist on a spectrum and therefore there will be individuals who could have opposite coded stances by even two human qualitative coders. Since the researcher is making a qualitative decision of when a person’s stance moves between two ends, the middle ground can often be coded either way. For example, is a person who owns guns for hunting but pushes for very strong hand gun control measures, pro gun or anti gun? Does wanting strong background checks or other restrictions inherently make someone pro/anti gun? Or does someone who has a strong stance against their personal gun ownership but also states they do not think background checks are necessary or valuable (i.e. guns aren’t for me but I don’t care about others), pro or anti gun control? Even in the context of a large body of tweets, it may not be fully clear how this stance should be coded and is not necessarily an error on either side but rather a decision reflecting this balance. An example is user 66811907 who has a large body of tweets that different coders could code differently “This your friendly reminder that New York already bans ”assault weapons,” that the shooter procured his NY-compliant ”non-assault weapon” legally, and that the cops who stopped him showed up with their ”assault weapon” version of the same gun to combat a civilian threat. https://t.co/KLmqwQM73f” and “Rifles of any kind account for fewer annual homicides than hands and feet. Mass shooters don’t need pistol grips, etc., to cause horrific carnage, but on the flip side those features can be quite useful for lawful owners using them in self-defense. That’s why cops carry them, too.” This example is just one that highlights that neither AI nor human coding is without issues, error, or disagreement. This is seen in the reporting of inter-rater reliability in qualitative coding and could be a useful measure to add to big data analysis.
The effective coding of large data sets with an aim toward integrating these two methodologies should be the goal of research aimed at informing the larger research field and policy. Without the purposeful integration of these two methodologies, much of this research, and its impact, will continue to be limited. Both guns and climate change have a range of historical contexts, political connotations and expectations, intersections with race and gender, and sometimes conspiracy theories that can play a role in how people talk about these ideas online and inform their stances. It is in this space that the subject matter expertise can provide that contextual information that can inform the work of AI algorithms. This concept has been discussed through the idea of grounding AI with use-case specific information that is outside of the training data. The social science insights provide the ground truth necessary to properly ground AI focused on complex human stances to ensure that the AI model is acknowledging the real-world context, to minimizing errors and maximizing the relevance of the generated outputs.
V Conclusion

### v2

@@ IV Results
We report the macro averaged precision, recall, and F1 score for each model, and we use the F1 score as the primary metric due to the class imbalance in the datasets. Metrics reported are averaged across five trials for all models.
IV-A Gun Control Dataset
TABLE IV: Results for stance labeling classification on the Gun Control dataset ($N=350$). The best F1 (macro) scores for each model type are shown in bold and the best overall scores are underlined.
@@ IV-A Gun Control Dataset
BERT-base-uncased 84.25 85.38 84.66 83.08 82.75 82.88
BERT-large-uncased 80.47 82.16 79.96 81.44 81.71 81.13
RoBERTa-base 83.33 83.94 83.60 84.79 83.99 84.25
RoBERTa-large 81.39 82.69 79.52 84.41 84.18 83.14
BERTweet 83.59 85.43 82.97 85.45 85.76 85.13
GPT-4 - - - 89.81 85.17 87.05
GraphSAGE 87.58 88.75 88.02 91.09 90.46 90.70
GAT 87.17 88.99 87.56 91.28 91.46 91.36
The classification results for the stance labeling models on the gun control dataset are shown in Table IV. On the gun control dataset, BERT and RoBERTa transformer methods achieve a macro averaged F1 score of between 82%-84% when evaluated on the set of manually annotated users. BERTweet, having been finetuned on tweets, surpasses the performance of the other transformer models achieving an F1 score of 85.13%. GPT-4 zero-shot classification outperforms the trained BERTweet classifier by two percent on F1 score and a four percent increase in precision while maintaining an equivalent recall performance.
For the gun control dataset, the GNN-based models that incorporate network structure information with the content-based representations of users produced by BERTweet, outperform both the trained text-based classifiers and zero-shot classification with GPT-4. Using graphSAGE and GAT to enrich users’ content-based representations results in an improvement in stance labeling classification compared with the unenriched BERTweet embeddings. With GAT, incorporating users’ local neighborhood information with their text-based embeddings improves classification performance by four percent in F1 score. Compared with zero-shot classification using GPT-4, the BERTweet + GAT model performs 3%-4% better in F1 score on classifying user stance labels for the manually labeled user set. Compared with graphSAGE, which takes the mean representation of a node’s neighborhood effectively weighting each incident edge on the node equally, GAT utilizes edge softmax attention to dynamically compute edge weights when aggregating node neighborhood information. The dynamic computation of edge importance for users results in a marginal improvement in classification performance.
IV-B Climate Change Dataset
TABLE V: Results for stance labeling classification on the Climate Change dataset ($N=250$). The best F1 (macro) scores for each model type are shown in bold and the best overall scores are underlined.
@@ IV-B Climate Change Dataset
BERT-base-uncased 76.74 79.82 77.36 78.60 82.54 79.36
BERT-large-uncased 75.63 79.14 73.86 76.11 80.28 74.17
RoBERTa-base 76.30 79.92 76.21 76.63 80.92 76.50
RoBERTa-large 75.72 79.36 75.16 76.91 81.33 76.29
BERTweet 77.26 80.78 77.64 78.35 82.66 78.78
GPT-4 - - - 85.36 86.08 85.71
GraphSAGE 84.62 86.06 85.25 83.80 85.91 84.67
GAT 83.73 86.43 84.70 83.47 87.02 84.62
The classification results for the stance labeling models on the climate change dataset are shown in Table V. Compared with the gun control dataset, where pro gun control users outnumber anti gun control users by roughly a ratio of 2:1 in the heuristic labeled user set, the imbalance in the climate change dataset is significant with climate change believers outnumbering disbelievers by a ratio of nearly 10:1. We also observe that climate change related conversations are less topic-focused than gun control discussions on Twitter. Due in part to the severe class imbalance in the dataset in addition to the less focused of climate change conversation activity on Twitter, stance determination of users in the climate change dataset is more difficult than in the gun control dataset.
Of the BERT and RoBERTa based classifier models, BERT-base-uncased performs the best with an F1 score of 79.36%, narrowly outperforming BERTweet with an F1 score of 78.78%. Enriching the BERTweet user embeddings with structural information from the user interaction graph improves the classification performance by six percent to 84.67% F1 score. Despite the performance uplift from the user social network, for the climate change dataset zero-shot classification with GPT-4 outperforms all other models with an F1 score of 85.71%, outperforming other transformer methods by six percent and GNN-based approaches by one percent.
In all, enriching text-based representations of users with structural information from the user-interaction graph results in a significant improvement in classification performance compared with the text-based embeddings alone. Zero-shot classification of user stances using GPT-4 consistently outperforms BERT and RoBERTa-based classifier models, but incorporating social network information surpasses GPT-4 classification in the gun control dataset and significantly narrows the performance gap in the climate change dataset.
V Discussion
@@
When researching online polarization, there are serious divides in how questions regarding user stance are structured, and this divide is also reflected in how different researchers integrate the literature and findings into future work. Oftentimes, this split can be seen in how social science researchers integrate computer science research into their studies and vice versa. This methodological gap can inhibit the understanding of the phenomena of online polarization and radicalization. Big data research can provide scalability not feasible by traditional qualitative social science research, while social science expertise provides understanding of the subtleties of these interactions on human behavior. As has been seen in the aftermath of events like radicalized mass shooters, the online behavior of these individuals is not well understood and the need to understand how and why these processes take place online needs to be examined. The integration of subject-matter expertise with AI is necessary for the large-scale application of big data methodology into human-level polarization research and policy work, requiring a balance of the nuance provided by social science subject-matter experts with the speed and volume provided by computation experts. Qualitative research offers methodological tools like narrative analysis to analyze the cultural and historical contingency of the terms, beliefs and issues narrators address in their writings [27]. The small narratives created from tweets can provide insights into the range of ongoing societal arguments that create collective identity online [28]. The wealth of this data adds critical information to understanding how people discuss their thoughts and opinions about polarizing topics like gun rights or climate change, and the ability to combine these two approaches can greatly benefit the understanding of online polarization.
@@
One area of difficulty in automating stance detection on topics like gun control or climate change is the need to dichotomize beliefs that really exist on a spectrum. The extreme ends of these beliefs are straight-forward to categorize, but the majority of people exist in the space in-between. A better goal is to qualitatively code the tipping point as people’s stance moves between stances on issues. Automated stance analysis can mislabel users because of difficulties in discerning written cues like tone, connotation, coded words, or pictures. Some of these coding errors are due to a purposeful misdirection of a tweet that is clarified by the picture. For example, user 2945287090 said “I’m pro-choice…are you? https://t.co/WtYW6zohSB” but the picture has them wearing a shirt that said pro-choice with an array of guns to “choose” from. Since the picture cannot be assessed by many stance labeling algorithms, these misdirections would be potentially miscoded. There are also users who frame their arguments in neutral language or the language of their opposition. For example, user 66811907 states “Gun control advocates do not have a monopoly on outrage or sorrow. They do not have a monopoly on ideas for addressing gun violence. They do not have a monopoly wanting to fix the problem.” The larger context of this user’s tweets balance between being pro-Second Amendment and being critical to the responses to mass shooting events, such as Uvalde. As a whole, this user falls more strongly on the side of being anti-gun control but certainly has criticism of open market gun purchasing. The neutral framing can be difficult to for automated approaches to parse since the full narrative of the users’ tweets need to be taken into account.
@@
However, the coding errors are not only on the side of computational approaches. One of the key limitations to using only human-based qualitative coding is that the deep understanding of stances is highly time-consuming and delayed. Especially in online space, the posting moves far faster than researchers can code by hand. This issue means that oftentimes qualitative researchers are forced to either focus on a subset of tweets or a subset of users, and when examining a subset of tweets to establish stance, there are bound to be errors. The balance between efficiency and accuracy means that there will be miscoded information that could potentially be corrected by looking at all the available data. The ability to integrate big data and computational methods can be a force multiplier for qualitative researchers as a mixed methods approach greatly improves the scope and generalizability of work on online polarization.
@@
As mentioned previously, these stances exist on a spectrum and since the researcher is making a qualitative decision of when a person’s stance moves between two ends, the middle ground can often be coded either way. For example, is a person who owns guns for hunting but pushes for very strong hand gun control measures, pro-gun control or anti-gun control? Does wanting strong background checks or other restrictions inherently make someone pro/anti gun? Even in the context of a large body of tweets, it may not be fully clear how this stance should be coded and is not necessarily an error on either side but rather a decision reflecting this balance. One example is user 66811907 who has a large body of tweets that different coders could code differently “This is your friendly reminder that New York already bans “assault weapons”, that the shooter procured his NY-compliant “non-assault weapon” legally, and that the cops who stopped him showed up with their “assault weapon” version of the same gun to combat a civilian threat.” and “Rifles of any kind account for fewer annual homicides than hands and feet. Mass shooters don’t need pistol grips, etc., to cause horrific carnage, but on the flip side those features can be quite useful for lawful owners using them in self-defense. That’s why cops carry them, too.” This example is just one that highlights that neither AI nor human coding is without issues, error, or disagreement and the reporting of inter-rater reliability could be a useful measure to add for both human annotators and automated approaches.
VI Conclusion
In this work, we take a step towards a broader understanding of user stance labeling through socially infused text mining—incorporating both textual content and network interactions using language models and graph neural networks. Both gun control and climate change are issues that have a range of historical contexts, political connotations, and intersections with race and gender. It is in this space that subject-matter expertise can provide the contextual information to inform the work of computational approaches. Social science insights provide the ground truths necessary to properly contextualize computational approaches to complex human stances to ensure that the models are acknowledging the real-world context, minimizing errors, and maximizing the relevance of the generated outputs to policymakers and to the broader research community.